OpenStudy (anonymous):

Anyone have a formula for this? Christian obtains a 25/9 balloon mortgage to finance $327,750 at 4.6%. What is the remaining balance when the balloon payment is due? a. $181,460.52 b.$198,762.12 c. $249,797.14 d. $271,180.37

OpenStudy (anonymous):

P= $327,750 F= (P+ i)^t i= 4.6/100 = 0.46 F= ($327,750 + 0.46)^t the number of years when this is due is not specified in the question